# An Efficient Genus Algorithm Based on Graph Rotations

### v2

@@ Abstract.
We study the problem of determining the minimal genus of a given finite connected graph. We present an algorithm which, for an arbitrary graph $G$ with $n$ vertices, determines the orientable genus of $G$ in $\mathcal{O}({2^{(n^{2}+3n)}}/{n^{(n+1)}})$ steps. This algorithm avoids the difficulties that many other genus algorithms have with handling bridge placements which is a well-known issue [32]. The algorithm has a number of properties that make it useful for practical use: it is simple to implement, it outputs the faces of the optimal embedding, it outputs a proof certificate for verification and it can be used to obtain upper and lower bounds. We illustrate the algorithm by determining the genus of the $(3,12)$ cage (which is 17); other graphs are also considered.
@@ 1.2. Main Result
PAGE described in §4 determines the genus of a connected multigraph $G(V,E)$ with runtime of $\mathcal{O}({2^{(n^{2}+3n)}}/{n^{(n+1)}})$.
@@
In 1963, Youngs established the fundamental principle of graph rotation theory, demonstrating that any embedding of a connected graph can be fully determined by the rotation of edges at its vertices [51]. For a vertex of degree $k$ there are $(k-1)!$ distinct rotations of edges at that vertex and thus for a $k$-regular graph of $n$ vertices there are $(k-1)!^{n}$ total embeddings into an orientable surface. It has been well established that the problem of determining the genus is NP-hard [47]. However, it is tractable for fixed genus [30, 32, 41]. Mohar has proven the existence of a linear time algorithm for arbitrary fixed surfaces [30]. However, the details to actually implement the algorithm have evaded the field for many decades. Similar efficient algorithms are more of theoretical interest than practical use given their large constant run-time factors, super-exponential scaling factors when genus is not kept fixed, and immense complexity that has prevented real-world implementation. The forbidden minor approach by Robertson and Seymour [41], for instance, is theoretically cubic time because it checks if a given graph has one of the finitely many forbidden minors for the fixed surface [42]. However, the complete list of forbidden minors is only known for planar [25] and projected plane graphs [1]. Even small toroidal graphs (genus 1) cannot be solved with this approach since there are at least 17.5K toroidal minors and likely many more [34]. In practice, the finite number of graph minors scales at an impractical super-exponential rate with the genus. This makes it intractable to compute all the minors (months of super-compute has been thrown at it to no avail). As it stands, the best implemented algorithms are exponential time [32, 7, 34] and the rest are too complex/impractical for implementation and have intractable constant factors.
Recently, an algorithm called multi_genus by Gunnar Brinkmann has emerged as a particularly fast method for computing genus on graphs with relatively low genus compared to the vertex degree [6]. It outperforms many previous algorithms, including ours, for graphs where vertex degrees exceed 5, scaling more effectively with edge count and vertex degree. However, our algorithm remains advantageous for graphs with vertices of degree 5 or lower. Additionally, our approach is comparatively simpler to implement and scales more efficiently with the genus, offering an advantageous alternative in cases where the genus is large. As it stands, multi_genus represents the fastest known approach for high-degree graphs, and low relative genus, whereas PAGE provides an effective alternative, and is particularly advantageous for graphs with bounded vertex degree.
@@ 2.2. Cages
Another graph family of special interest is the $(r,g)$-cage graphs. They are the smallest $r$-regular graphs with girth $g$. The genus of $(3,g)$ cage graphs is known up to $g=10$, and PAGE extends these results by determining the genus of the $(3,12)$ cage. Although the structure of $(3,g)$ cages is not fully known for $g>12$, our approach is likely applicable to higher values of $g$ as these cages are discovered. We outperform all existing algorithms, including multi_genus, for $g>8$, and have the only tractable algorithm for $g\geq 12$.
@@ 3. Progressive Refinement of Genus Bounds
$\frac{m-4n/3+2}{2}=g_{0}\leq g_{1}\leq\cdots\leq$ $g\leq\cdots G_{1}\leq G_{0}=\frac{m-n+2}{2}$
@@
with $g_{k},G_{k}$ computed in runtime $\mathcal{O}(((n-1)!/k/(n-k)!)^{n/3-k})$ for $k>0$ and in constant time using the formulas for $k=0$. In other words, the sequence of lower bounds $g_{k}$ and sequence of upper bounds $G_{k}$ converge to the genus $g$ after $\ell/6$ iterations with intermediate absolute error bounded by $\ell/6-k$.
@@ 4. Proof of the Algorithm
We construct PAGE with formal pseudocode. The general idea is Euler’s formula $n-m+F=2-2g$ which links the facial walks $F$ with the genus $g$ [14]. Naively, finding the minimum genus amounts to searching all combinations of cycles to find the one that corresponds to a maximal fitting of facial walks. Traditional improved exhaustive search algorithms instead search through the rotation systems since they each induce a valid facial walk whereas not all cycle combinations are valid. Searching through rotation systems however does not easily allow further pruning of the search space nor inform a heuristic search through rotation systems in an order that most quickly narrows down the genus. Our algorithm instead searches through all cycle combinations which facilitates a number of optimizations (early stopping, heuristic search) and, with our main contribution, still allows us to prune invalid rotation systems which results in an exponentially reduced search space.
@@
Given a graph $G$ with finite $n$ vertices and finite $m$ undirected edges, the set of simple cycles $C$ of $G$ is finite and any valid set of facial walks of $G$ is a subset of $C$.
@@
The finite size of $C$ follows from the finite combinations of up to $n$ vertices and that a simple cycle cannot repeat a vertex and must therefore contain at most $n$ vertices. Since each facial walk in a set of valid facial walks of $G$ is a simple cycle, the lemma is clear. ∎
By Lemma 6, we need to filter out a subset of the simple cycles. Finding all the simple cycles of a graph, can be done in $O((c+1)(n+m))$ where $c$ is the number of cycles using Johnson’s Algorithm [20]. Since we do not need to generate all the cycles, only the ones of a few lengths, the algorithm by Liu et al. [26] is advantageous for its simplicity, easily parallelized form, and ability to generate all cycles of a given length $k$ efficiently without keeping other cycles in memory or doing extensive computation on the graph beforehand.
Algorithm 1 $k$-Cycle Finding Algorithm
@@
2:  Output: Sequence of length-$k$ cycles of $G$
@@
9:     else
10:        for each neighbor $v$ of last vertex in path do
11:           if $v\notin\text{path}$ and $v>\text{first vertex of path}$ then
12:              enqueue(queue, path)
@@
14:        end for
15:     end if
16:  end while
@@
Given a graph $G$ with finite $n$ vertices and finite $m$ undirected edges, choose any vertex $v\in V$ of degree $d$ and all valid set of facial walks must include $v$ exactly $d$ times.
@@
A valid set of facial walks must use each directed edge of $G$ exactly once. Since each facial walk is a cycle, it must use 2 of the directed edges that touch $v$ or none of them. There are $2\cdot d$ directed edges that touch $v$ and thus $2\cdot d/2=d$ facial walks including $v$. ∎
By Lemma 7, we can annotate our adjacency list with the number of uses of each vertex and use that in a PotentialMaxFit procedure to reject cycle combinations early if they try to use a vertex more than its degree number of times.
@@
Given a graph $G$ with finite $n$ vertices and finite $m$ undirected edges, any valid set of facial walks must be a set of simple cycles whose lengths add up to the number of directed edges $2\cdot m$.
@@
The length of a simple cycle is the number of unique edges it uses. A valid set of facial walks must by definition use each directed edge of $G$ exactly once. ∎
An important way to prune the search space is to not blindly try all combinations of cycles and instead only try valid cycle combinations that could add up to use all the directed edges. Checking all valid cycle distributions to find the working one with the most cycles would then yield the genus as formalized in Lemma 8. In order to stop early when a working distribution is found, it is desired to iterate in order of decreasing number of cycles. Finding the cycle distribution with the most cycles is the bounded integer knapsack problem in disguise (each cycle has value 1, costs its length, and there is a bounded number of cycles of each length) so it is NP-hard by reduction. Various pseudo-polynomial solutions for this problem exist. In practice, iterating in sorted order induces too much of an overhead. We instead include pseudocode for iterating in increasing order of max cycle length. This also saves the overhead of computing all cycle lengths. It is trivial to sort the produced cycle distributions in linear time (linear in the number of cycle distributions) using bucket sort.
@@
Given a graph $G$ with finite $n$ vertices and finite $m$ undirected edges, a valid set of facial walks is a set of simple cycles that use each directed edge exactly once and do not contain two cycles $c_{1},c_{2}$ such that $c_{1}$ contains adjacent directed edges $(a,b),(b,c)$ and $c_{2}$ contains $(c,b),(b,a)$ for any vertex $b$ with degree greater than $2$.
@@
A facial walk must be a simple cycle. A valid set of facial walks, is 11 a valid rotation system. A valid rotation system is a permutation of the incident edges to each vertex determining that entering through one edge requires leaving through the next. A facial walk can never contain both edges $(u,v)$ and $(v,u)$ so $a\neq b$. If $c_{1}$ and $c_{2}$ exist, the degree of $b$ must be $2$ for the two vertices to be next to each other when reading the permutation in either order which is a contradiction. ∎
@@
The idea of the algorithm is to go through each cycle distribution (discarding ones that are not relevant to finding the maximum cycle distribution size, and thus the genus, because a cycle distribution with more cycles has already been found to work). For each cycle distribution, the algorithm then recursively picks a cycle until either the cycle distribution is fully used (thus all directed edges are used exactly once, and the distribution works) or any cycle length does not have cycles with unused edges available (thus the distribution could never use all unused edges exactly once, and does not work). Another optimization the algorithm makes is choosing cycles from the same vertex until that vertex is satisfied (one cycle for each incident un-directed edge used) since that is a necessary condition to all edges being used exactly once and allows failing early as soon as a vertex cannot be satisfied. The recursion depth is small since the maximum number of cycles that can fit is $O(m-n)$ and in most cases the algorithm will fail out of the recursion early.
@@
PAGE is $\mathcal{O}(2^{n^{2}+3}/n^{n+1})$ for complete graphs.
@@
    Finding all c elementary cycles of a graph: $\mathcal{O}((c+1)(n+m))=\mathcal{O}(cn+cm)$ using Johnson’s algorithm.
@@
    Organizing by vertex: $\mathcal{O}(cn)$ by iterating through the cycles of length at most $n$.
@@
    Iterating through all cycles of length $k$: $\mathcal{O}(n!/k/(n-k)!)$ since this is the number of $k$-length cycles in a complete graph. This is worst case $\mathcal{O}(n!/((n/2)!)^{2})$ which is roughly $\mathcal{O}(2^{n})$.
@@
    Find the most used vertex to explore: $\mathcal{O}(n)$ by iterating through the vertices (overall time complexity is better when write is kept $\mathcal{O}(1)$).
@@
    Looking up the cycles that use a vertex: $\mathcal{O}(1)$ via hashset lookup.
@@
    Checking if a cycle is used: $\mathcal{O}(1)$ via hashset lookup.
@@
    Checking if the edges of a cycle are used: $\mathcal{O}(e)$ where e is the number of edges in the cycle by storing the edges used in a hashset.
@@
    Checking $ijk$ criterion of a cycle: $\mathcal{O}(e)$ by storing the current rotation with the adjacency list.
@@
    Search iteration ($f=$ implied fit, $b=$ cycles by vertex $\geq u=$ unused $\geq a=$ w/ edges available $\geq d=$ ijk good): $T(f)=\mathcal{O}(n)+\mathcal{O}(1)+b\mathcal{O}(1)+u\mathcal{O}(e)+a\mathcal{O}(e)+dT(f-1);T(0)=0\implies\mathcal{O}((d^{f}\cdot(n+b+ue+ae))=\mathcal{O}(d^{f}\cdot(n+b+e(u+a)))<\mathcal{O}((b^{f+2})$.
@@
    All search iterations ($t=$number of start cycles to try out $<n2^{n}$): $\mathcal{O}(b^{f+2}\cdot t)<\mathcal{O}((2^{n}/n)^{n+2}\cdot n2^{n})=\mathcal{O}(2^{n^{2}+3n}/n^{n+1})$.

### v3

@@ Abstract.
We study the problem of determining the minimal genus of a simple finite connected graph. We present an algorithm which, for an arbitrary graph $G$ with $n$ vertices, determines the orientable genus of $G$ in $\mathcal{O}(n(4^{m}/n)^{n/t})$ steps where $t$ is the girth. This algorithm avoids the difficulties that many other genus algorithms have with handling bridge placements which is a well-known issue [32]. The algorithm has a number of properties that make it useful for practical use: it is simple to implement, it outputs the faces of the optimal embedding, it outputs a proof certificate for verification and it can be used to obtain upper and lower bounds. We illustrate the algorithm by determining the genus of the $(3,12)$ cage (which is 17); other graphs are also considered.
@@ 1.2. Main Result
PAGE described in §4 determines the genus of a connected multigraph $G(V,E)$ with runtime of $\mathcal{O}(n(4^{m}/n)^{n/t})$ where $t$ is the girth.
@@
In 1963, Youngs established the fundamental principle of graph rotation theory, demonstrating that any embedding of a connected graph can be fully determined by the rotation of edges at its vertices [51]. For a vertex of degree $k$ there are $(k-1)!$ distinct rotations of edges at that vertex and thus for a $k$-regular graph of $n$ vertices there are $(k-1)!^{n}$ total embeddings into an orientable surface. It has been well established that the problem of determining the genus is NP-hard [47]. However, it is tractable for fixed genus [30, 32, 41]. Mohar has proven the existence of a linear time algorithm for arbitrary fixed surfaces [30]. However, the details to actually implement the algorithm have evaded the field for many decades. Similar efficient algorithms are more of theoretical interest than practical use given their large constant run-time factors, super-exponential scaling factors when genus is not kept fixed, and immense complexity that has prevented real-world implementation. The forbidden minor approach by Robertson and Seymour [41], for instance, is theoretically cubic time because it checks if a given graph has one of the finitely many forbidden minors for the fixed surface [42]. However, the complete list of forbidden minors is only known for planar [25] and projective plane graphs [1]. Even small toroidal graphs (genus 1) cannot be solved with this approach since there are at least 17.5K toroidal minors and likely many more [34]. In practice, the finite number of graph minors scales at an impractical super-exponential rate with the genus. This makes it intractable to compute all the minors (months of super-compute has been thrown at it to no avail). As it stands, the best implemented algorithms are exponential time [32, 7, 34] and the rest are too complex/impractical for implementation and have intractable constant factors.
Recently, an algorithm called multi_genus by Gunnar Brinkmann has emerged as a particularly fast method for computing genus on graphs with relatively low genus compared to the vertex degree [6]. It outperforms many previous algorithms, including ours, for graphs where vertex degrees exceed 5, scaling more effectively with edge count and vertex degree. However, our algorithm remains advantageous for graphs with vertices of degree 5 or lower. Additionally, our approach is comparatively simpler to implement and scales more efficiently with the genus, offering an advantageous alternative in cases where the genus is large. As it stands, multi_genus represents the fastest known approach for high-degree graphs, and low relative genus, whereas PAGE provides an effective alternative, and is particularly advantageous for graphs with bounded vertex degree or high girth.
@@ 2.2. Cages
Another graph family of special interest is the $(r,t)$-cage graphs. They are the smallest $r$-regular graphs with girth $t$. The genus of $(3,t)$ cage graphs is known up to $t=10$, and PAGE extends these results by determining the genus of the $(3,12)$ cage. Although the structure of $(3,t)$ cages is not fully known for $t>12$, our approach is likely applicable to higher values of $t$ as these cages are discovered. We outperform all existing algorithms, including multi_genus, for $t>8$, and have the only tractable algorithm for $t\geq 12$.
@@ 3. Progressive Refinement of Genus Bounds
$\frac{m-4n/3+2}{2}=g_{0}\leq g_{1}\leq\cdots\leq$ $g\leq\cdots\leq G_{1}\leq G_{0}=\frac{m-n+2}{2}$
@@
with $g_{k},G_{k}$ computed in runtime $\mathcal{O}(((2m)!/k!/(2m-k)!)^{n/3-k})$ for $k>0$ and in constant time using the formulas for $k=0$. In other words, the sequence of lower bounds $g_{k}$ and sequence of upper bounds $G_{k}$ converge to the genus $g$ after $\ell/6$ iterations with intermediate absolute error bounded by $\ell/6-k$.
@@ 4. Proof of the Algorithm
We construct PAGE with formal pseudocode. The general idea is Euler’s formula $n-m+F=2-2g$ which links the facial walks $F$ with the genus $g$ [14]. Naively, finding the minimum genus amounts to searching all combinations of closed directed trails to find the one that corresponds to a maximal fitting of facial walks. Traditional improved exhaustive search algorithms instead search through the rotation systems since they each induce a valid facial walk whereas not all trail combinations are valid. Searching through rotation systems however does not easily allow further pruning of the search space nor inform a heuristic search through rotation systems in an order that most quickly narrows down the genus. Our algorithm instead searches through all non-backtracking closed directed trail combinations which facilitates a number of optimizations (early stopping, heuristic search) and, with our main contribution, still allows us to prune invalid rotation systems which results in an exponentially reduced search space.
@@
Given a graph $G$ with finite $n$ vertices and finite $m$ undirected edges, the set of non-backtracking closed trails $C$ of $G$ is finite and any valid set of facial walks of $G$ is a subset of $C$.
@@
The finite size of $C$ follows from the finite combinations of up to $2m$ directed edges and that a trail cannot repeat an edge and must therefore contain at most $2m$ edges. Since each facial walk in a set of valid facial walks of $G$ is a non-backtracking closed trail, the lemma is clear. ∎
By Lemma 6, we need to filter out a subset of the non-backtracking closed directed trails. Since we do not need to generate all the trails, only the ones of a few lengths, we adapt the algorithm by Liu et al. [26] for enumerating simple cycles. It is advantageous for its simplicity, easily parallelized form, and ability to generate all cycles of a given length $k$ efficiently without keeping other cycles in memory or doing extensive computation on the graph beforehand. It is easily extended to non-backtracking closed directed trails as seen in the below pseudo-code:
Algorithm 1 $k$-trail Finding Algorithm
@@
2:  Output: Sequence of length-$k$ non-backtracking closed directed trails of $G$
@@
8:        if the first vertex of path is a neighbor of the last vertex then
9:           if the last vertex is not the same as the second vertex then
10:              if if the directed edge from the last vertex to the first vertex is not a repeat then
@@
12:              end if
@@
14:        end if
15:     else
16:        for each neighbor $v$ of last vertex in path do
[…]
17:              continue
18:           end if
19:           if the directed edge from the last vertex to $v$ is a repeat then
19:              continue
20:           end if
21:           if $v\geq\text{first vertex of path}$ then
22:              enqueue(queue, path)
23:           end if
24:        end for
25:     end if
26:  end while
@@
Given a graph $G$ with finite $n$ vertices and finite $m$ undirected edges, choose any vertex $v\in V$ of degree $d$ and all valid sets of facial walks must include $v$ exactly $d$ times.
@@
A valid set of facial walks must use each directed edge of $G$ exactly once. Since each facial walk is a directed trail, it must use 2 of the directed edges that touch $v$ or none of them. There are $2\cdot d$ directed edges that touch $v$ and thus $2\cdot d/2=d$ facial walks including $v$. ∎
By Lemma 7, we can annotate our adjacency list with the number of uses of each vertex and use that in a PotentialMaxFit procedure to reject trail combinations early if they try to use a vertex more than its degree number of times.
@@
Given a graph $G$ with finite $n$ vertices and finite $m$ undirected edges, any valid set of facial walks must be a set of non-backtracking closed directed trails whose lengths add up to the number of directed edges $2\cdot m$.
@@
The length of a trail is the number of unique edges it uses. A valid set of facial walks must by definition use each directed edge of $G$ exactly once. ∎
An important way to prune the search space is to not blindly try all combinations of trails and instead only try valid trail combinations that could add up to use all the directed edges. Checking all valid cycle distributions to find the working one with the most cycles would then yield the genus as formalized in Lemma 8. In order to stop early when a working distribution is found, it is desired to iterate in order of decreasing number of cycles. Finding the cycle distribution with the most cycles is the bounded integer knapsack problem in disguise (each cycle has value 1, costs its length, and there is a bounded number of cycles of each length) so it is NP-hard by reduction. Various pseudo-polynomial solutions for this problem exist. In practice, iterating in sorted order induces too much of an overhead. We instead include pseudo-code for iterating in increasing order of max cycle length. This also saves the overhead of computing all cycle lengths. It is trivial to sort the produced cycle distributions in linear time (linear in the number of cycle distributions) using bucket sort.
@@
Given a graph $G$ with finite $n$ vertices and finite $m$ undirected edges, a valid set of facial walks is a set of non-backtracking closed directed trails that use each directed edge exactly once and do not contain two trails $c_{1},c_{2}$ such that $c_{1}$ contains adjacent directed edges $(a,b),(b,c)$ and $c_{2}$ contains $(c,b),(b,a)$ for any vertex $b$ with degree greater than $2$.
@@
A facial walk must be a non-backtracking closed directed trail. A valid set of facial walks, is 11 a valid rotation system. A valid rotation system is a permutation of the incident edges to each vertex determining that entering through one edge requires leaving through the next. A facial walk can never contain both edges $(u,v)$ and $(v,u)$ so $a\neq b$. If $c_{1}$ and $c_{2}$ exist, the degree of $b$ must be $2$ for the two vertices to be next to each other when reading the permutation in either order which is a contradiction. ∎
@@
The idea of the algorithm is to go through each trail distribution (discarding ones that are not relevant to finding the maximum trail distribution size, and thus the genus, because a trail distribution with more trails has already been found to work). For each trail distribution, the algorithm then recursively picks a trail until either the trail distribution is fully used (thus all directed edges are used exactly once, and the distribution works) or any trail length does not have trails with unused edges available (thus the distribution could never use all unused edges exactly once, and does not work). Another optimization the algorithm makes is choosing trails from the same vertex until that vertex is satisfied (one trail for each incident un-directed edge used) since that is a necessary condition to all edges being used exactly once and allows failing early as soon as a vertex cannot be satisfied. The recursion depth is small since the maximum number of cycles that can fit is $O(m-n)$ and in most cases the algorithm will fail out of the recursion early.
@@
PAGE is $\mathcal{O}(n(4^{m}/n)^{n/t})$ for graphs of girth $t$.
@@
    The number of directed trails of any length is $\mathcal{O}(4^{m})$ because it is the sum of the values in the $2m$-th row of Pascal’s triangle.
@@
    This can also be used to upper bound the number of non-backtracking closed directed trails which is what our trail finding algorithm enumerates. Let $c$ be the number of trails enumerated by our trail finding algorithm.
@@
    Organizing the $c$ trails by vertex is $\mathcal{O}(2mc)<\mathcal{O}(2m4^{m})$ since each trail has at most $2m$ edges.
@@
    Finding the most used vertex to explore: $O(n)$ by iterating through the vertices (overall time complexity is better when write is kept $O(1)$)
@@
    Looking up the cycles that use a vertex: $O(1)$ via hashset lookup
@@
    Checking if a cycle is used: $O(1)$ via hashset lookup
@@
    Checking if the edges of a cycle are used: $O(m)$ by storing the edges used in a hashset
@@
    Checking if adding cycle breaks rotation system: $O(m)$ by storing the current rotation with the adjacency list
@@
    Search iteration ($f=$ implied fit, $b=$ cycles by vertex $\geq u=$ unused $\geq a=$ w/ edges available $\geq d=$ rotation valid): $T(f)=\mathcal{O}(n)+\mathcal{O}(1)+b\mathcal{O}(1)+u\mathcal{O}(m)+a\mathcal{O}(m)+dT(f-1);T(0)=0\implies\mathcal{O}(d^{f}\cdot(n+b+um+am))=\mathcal{O}(d^{f}\cdot(n+b+m(u+a)))<\mathcal{O}(b^{f+2})$
@@
    All search iterations (h = number of start cycles to try out ¡ $4^{m}$): $\mathcal{O}(b^{f+2}\cdot h)<\mathcal{O}((4^{m}/n)^{n/t}\cdot 4^{m})=\mathcal{O}(n(4^{m}/n)^{n/t})$.